プリンキピア

第3章確率

3.3 Two Basic Rules of Probability3.3 確率の二つの基本法則

When calculating probability, there are two rules to consider when determining if two events are independent or dependent and if they are mutually exclusive or not.
確率を計算する際、二つの事象が独立か従属か、また排反かどうかを判断するときに、考慮すべき二つの法則がある。

The Multiplication Rule乗法定理

If A and B are two events defined on a sample space, then: P(A AND B) = P(B)P(A|B).
AとBが標本空間上で定義された二つの事象なら、P(A AND B) = P(B)P(A|B)となる。
This rule may also be written as: P(A|B) = P(A AND B) P(B)
この法則は、P(A|B) = P(A AND B) P(B) とも書ける。
(The probability of A given B equals the probability of A and B divided by the probability of B.)
(Bが起こったという条件のもとでのAの確率は、AとBの両方が起こる確率をBの確率で割ったものに等しい。)
If A and B are independent, then P(A|B) = P(A). Then P(A AND B) = P(A|B)P(B) becomes P(A AND B) = P(A)P(B).
AとBが独立なら、P(A|B) = P(A)である。したがって、P(A AND B) = P(A|B)P(B)は、P(A AND B) = P(A)P(B)となる。

The Addition Rule加法定理

If A and B are defined on a sample space, then: P(A OR B) = P(A) + P(B) - P(A AND B).
AとBが標本空間上で定義されているなら、P(A OR B) = P(A) + P(B) − P(A AND B)となる。
If A and B are mutually exclusive, then P(A AND B) = 0. Then P(A OR B) = P(A) + P(B) - P(A AND B) becomes P(A OR B) = P(A) + P(B).
AとBが排反なら、P(A AND B) = 0である。したがって、P(A OR B) = P(A) + P(B) − P(A AND B)は、P(A OR B) = P(A) + P(B)となる。
Example 3.14例題 3.14
Klaus is trying to choose where to go on vacation. His two choices are: A = New Zealand and B = Alaska
英語のヒント
クラウスは休暇でどこへ行くか選ぼうとしている。二つの選択肢は、A = ニュージーランド、B = アラスカである。
  • Klaus can only afford one vacation. The probability that he chooses A is P(A) = 0.6 and the probability that he chooses B is P(B) = 0.35.
    英語のヒント
    クラウスには休暇旅行1回分の費用しかない。Aを選ぶ確率はP(A) = 0.6、Bを選ぶ確率はP(B) = 0.35である。
  • P(A AND B) = 0 because Klaus can only afford to take one vacation
    クラウスが費用を払える休暇旅行は1回だけなので、P(A AND B) = 0である。
  • Therefore, the probability that he chooses either New Zealand or Alaska is P(A OR B) = P(A) + P(B) = 0.6 + 0.35 = 0.95. Note that the probability that he does not choose to go anywhere on vacation must be 0.05.
    したがって、ニュージーランドかアラスカのいずれかを選ぶ確率は、P(A OR B) = P(A) + P(B) = 0.6 + 0.35 = 0.95となる。休暇でどこにも行かないと決める確率は、0.05でなければならないことに注意せよ。
Example 3.15例題 3.15
Carlos plays college soccer. He makes a goal 65% of the time he shoots. Carlos is going to attempt two goals in a row in the next game. A = the event Carlos is successful on his first attempt. P(A) = 0.65. B = the event Carlos is successful on his second attempt. P(B) = 0.65. Carlos tends to shoot in streaks. The probability that he makes the second goal GIVEN that he made the first goal is 0.90.
英語のヒント
カルロスは大学でサッカーをしている。シュートを打つと65%の割合で得点する。次の試合では2回続けて得点を狙う予定である。A = カルロスが1回目に成功する事象、とすると、P(A) = 0.65である。B = 2回目に成功する事象、とすると、P(B) = 0.65である。カルロスには成功が続く傾向がある。1回目に得点したという条件のもとで、2回目にも得点する確率は0.90である。
a. What is the probability that he makes both goals?
a. 両方とも得点する確率はいくらか。
b. What is the probability that Carlos makes either the first goal or the second goal?
b. カルロスが1回目または2回目に得点する確率はいくらか。
c. Are A and B independent?
c. AとBは独立か。
d. Are A and B mutually exclusive?
d. AとBは排反か。
Solution解き方
a. The problem is asking you to find P(A AND B) = P(B AND A). Since P(B|A) = 0.90: P(B AND A) = P(B|A) P(A) = (0.90)(0.65) = 0.585
a. 求めるのはP(A AND B) = P(B AND A)である。P(B|A) = 0.90なので、P(B AND A) = P(B|A)P(A) = (0.90)(0.65) = 0.585となる。
Carlos makes the first and second goals with probability 0.585.
カルロスは0.585の確率で、1回目と2回目の両方で得点する。
b. The problem is asking you to find P(A OR B).
b. この問題で求めるのはP(A OR B)である。
P(A OR B) = P(A) + P(B) - P(A AND B) = 0.65 + 0.65 - 0.585 = 0.715
Carlos makes either the first goal or the second goal with probability 0.715.
カルロスは0.715の確率で、1回目または2回目に得点する。
c. No, they are not, because P(B AND A) = 0.585.
c. いいえ、独立ではない。P(B AND A) = 0.585だからである。
P(B)P(A) = (0.65)(0.65) = 0.423
0.423 ≠ 0.585 = P(B AND A)
So, P(B AND A) is not equal to P(B)P(A).
したがって、P(B AND A)はP(B)P(A)と等しくない。
d. No, they are not because P(A and B) = 0.585.
d. いいえ、排反ではない。P(A and B) = 0.585だからである。
To be mutually exclusive, P(A AND B) must equal zero.
排反であるためには、P(A AND B)が0でなければならない。
Example 3.16例題 3.16
A community swim team has 150 members. Seventy-five of the members are advanced swimmers. Forty-seven of the members are intermediate swimmers. The remainder are novice swimmers. Forty of the advanced swimmers practice four times a week. Thirty of the intermediate swimmers practice four times a week. Ten of the novice swimmers practice four times a week. Suppose one member of the swim team is chosen randomly.
英語のヒント
地域の水泳チームには150人が所属している。75人は上級者、47人は中級者、残りは初心者である。上級者の40人、中級者の30人、初心者の10人は週4回練習する。チームから1人を無作為に選ぶとする。
a. What is the probability that the member is a novice swimmer?
a. その人が初心者である確率はいくらか。
b. What is the probability that the member practices four times a week?
b. その人が週4回練習する確率はいくらか。
c. What is the probability that the member is an advanced swimmer and practices four times a week?
c. その人が上級者であり、かつ週4回練習する確率はいくらか。
d. What is the probability that a member is an advanced swimmer and an intermediate swimmer? Are being an advanced swimmer and an intermediate swimmer mutually exclusive? Why or why not?
d. その人が上級者であり、かつ中級者である確率はいくらか。上級者である事象と中級者である事象は排反か。その理由は何か。
e. Are being a novice swimmer and practicing four times a week independent events? Why or why not?
e. 初心者である事象と週4回練習する事象は独立か。その理由は何か。
Solution解き方
a. 28150
b. 80150
c. 40150
d. P(advanced AND intermediate) = 0, so these are mutually exclusive events. A swimmer cannot be an advanced swimmer and an intermediate swimmer at the same time.
d. P(上級者 AND 中級者) = 0なので、これらは排反な事象である。一人の泳者が同時に上級者かつ中級者になることはない。
e. No, these are not independent events.
P(novice AND practices four times per week) = 0.0667
P(novice)P(practices four times per week) = 0.0996
0.0667 ≠ 0.0996
e. いいえ、独立ではない。P(初心者 AND 週4回練習) = 0.0667、P(初心者)P(週4回練習) = 0.0996であり、0.0667 ≠ 0.0996である。
Example 3.17例題 3.17
Felicity attends Modesto JC in Modesto, CA. The probability that Felicity enrolls in a math class is 0.2 and the probability that she enrolls in a speech class is 0.65. The probability that she enrolls in a math class GIVEN that she enrolls in speech class is 0.25.
英語のヒント
フェリシティはカリフォルニア州モデストのModesto JCに通っている。数学の授業を履修する確率は0.2、スピーチの授業を履修する確率は0.65である。スピーチを履修するという条件のもとで、数学を履修する確率は0.25である。
Let: M = math class, S = speech class, M|S = math given speech
M = 数学の授業、S = スピーチの授業、M|S = スピーチを履修するという条件のもとで数学を履修する、とする。
  1. What is the probability that Felicity enrolls in math and speech?
    Find P(M AND S) = P(M|S)P(S).
    フェリシティが数学とスピーチの両方を履修する確率はいくらか。P(M AND S) = P(M|S)P(S)を求めよ。
  2. What is the probability that Felicity enrolls in math or speech classes?
    Find P(M OR S) = P(M) + P(S) - P(M AND S).
    フェリシティが数学またはスピーチの授業を履修する確率はいくらか。P(M OR S) = P(M) + P(S) − P(M AND S)を求めよ。
  3. Are M and S independent? Is P(M|S) = P(M)?
    MとSは独立か。P(M|S) = P(M)か。
  4. Are M and S mutually exclusive? Is P(M AND S) = 0?
    MとSは排反か。P(M AND S) = 0か。
Solution解き方
a. 0.1625, b. 0.6875, c. No, d. No
a. 0.1625、b. 0.6875、c. いいえ、d. いいえ。
Example 3.18例題 3.18
Studies show that about one woman in seven (approximately 14.3%) who live to be 90 will develop breast cancer. Suppose that of those women who develop breast cancer, a test is negative 2% of the time. Also suppose that in the general population of women, the test for breast cancer is negative about 85% of the time. Let B = woman develops breast cancer and let N = tests negative. Suppose one woman is selected at random.
英語のヒント
研究によれば、90歳まで生きる女性の約7人に1人(約14.3%)が乳がんを発症する。乳がんを発症する女性のうち、検査で陰性になる割合を2%とする。また、女性の一般母集団では、乳がん検査で陰性になる割合を約85%とする。B = 女性が乳がんを発症する、N = 検査で陰性となる、と置く。女性を1人無作為に選ぶとする。
a. What is the probability that the woman develops breast cancer? What is the probability that woman tests negative?
a. その女性が乳がんを発症する確率はいくらか。検査で陰性となる確率はいくらか。
b. Given that the woman has breast cancer, what is the probability that she tests negative?
b. その女性が乳がんであるという条件のもとで、検査で陰性となる確率はいくらか。
c. What is the probability that the woman has breast cancer AND tests negative?
c. その女性が乳がんであり、かつ検査で陰性となる確率はいくらか。
d. What is the probability that the woman has breast cancer or tests negative?
d. その女性が乳がんである、または検査で陰性となる確率はいくらか。
e. Are having breast cancer and testing negative independent events?
e. 乳がんである事象と検査で陰性となる事象は独立か。
f. Are having breast cancer and testing negative mutually exclusive?
f. 乳がんである事象と検査で陰性となる事象は排反か。
Solution解き方
a. P(B) = 0.143; P(N) = 0.85
b. P(N|B) = 0.02
c. P(B AND N) = P(B)P(N|B) = (0.143)(0.02) = 0.0029
d. P(B OR N) = P(B) + P(N) - P(B AND N) = 0.143 + 0.85 - 0.0029 = 0.9901
e. No. P(N) = 0.85; P(N|B) = 0.02. So, P(N|B) does not equal P(N).
e. いいえ。P(N) = 0.85、P(N|B) = 0.02なので、P(N|B)はP(N)に等しくない。
f. No. P(B AND N) = 0.0029. For B and N to be mutually exclusive, P(B AND N) must be zero.
f. いいえ。P(B AND N) = 0.0029である。BとNが排反であるためには、P(B AND N)が0でなければならない。
Example 3.19例題 3.19
Refer to the information in Example 3.18. P = tests positive.
例3.18の情報を参照せよ。P = 検査で陽性となる、とする。
  1. Given that a woman develops breast cancer, what is the probability that she tests positive. Find P(P|B) = 1 - P(N|B).
    女性が乳がんを発症するという条件のもとで、検査で陽性となる確率はいくらか。P(P|B) = 1 − P(N|B)を求めよ。
  2. What is the probability that a woman develops breast cancer and tests positive. Find P(B AND P) = P(P|B)P(B).
    女性が乳がんを発症し、かつ検査で陽性となる確率はいくらか。P(B AND P) = P(P|B)P(B)を求めよ。
  3. What is the probability that a woman does not develop breast cancer. Find P(B′) = 1 - P(B).
    女性が乳がんを発症しない確率はいくらか。P(B′) = 1 − P(B)を求めよ。
  4. What is the probability that a woman tests positive for breast cancer. Find P(P) = 1 - P(N).
    女性が乳がん検査で陽性となる確率はいくらか。P(P) = 1 − P(N)を求めよ。
Solution解き方
a. 0.98; b. 0.1401; c. 0.857; d. 0.15